OpenStudy (anonymous):

how to calculate the area of the circle r=10sintheta?

OpenStudy (anonymous):

the area of a circle = 2pi r^2 so you have r= 10 sin(theta)

OpenStudy (anonymous):

r^2=100 sin^2(theta)

OpenStudy (anonymous):

\[area = 200\pi \sin ^{2}(\theta)\]